a fixed perimeter of 164ft, what dimensions would yield the maximum area? what is the area?
Answer by stanbon(75887)

a fixed perimeter of 164ft,
what dimensions would yield the maximum area?
what is the area?
--------------------------------
2(L+W) = 164
L+W = 82
------------------
L = 82-W
--------------
Area = W(82-W) = 82W-W^2
maximum area occurs when W = -b/2a = -82/(-2) = 41
------------------
A(41) = 41*(82-41) = 1681 sq. ft. (maximum area)
===================================================
Cheers,
Stan H.
========================================